Operations Report – January 11th



Crew 272 Operations Report 11-1-2023

SOL: 10

Name of person filing report: Mason Kuhn

Non-nominal systems: Oven light, Suit 7’s vent panel on the back of the suit (one of the screw holes in the wood)

Notes on non-nominal systems: Noticed the oven light today while cooking dinner. It is definitely not an urgent problem. 1 of 2 screw holes in the wood to secure the vent panel to suit 7 is too large for the washer and nut. Ideally, the wood should be repaired/replaced, but I can temporarily secure the vent panel with a larger diameter washer and nut.


Spirit rover used: No

Hours: 207.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 111.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 219.0

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: None were used on the EVA to Hab Ridge.

Summary of Hab operations: Suit 7 repairs were made on the lower deck, see suits summary for information. 4 showers were taken today, so the water usage was above our average.

WATER USE: 26.1 gallons

Water (static tank): 180.8 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suit 7 has been repaired! There was a loose connection in terminal 3 of the toggle switch, and the wire-organizing zip tie was too tight. These items were resulting in spotty functionality. A new zip tie is still organizing the wires inside the suit back, but it is now a bit looser. The terminal 3 connection was retightened to the underside of the toggle switch, and Suit 7 electrical systems are now nominal. Suits 8,9,10,11 were used on the EVA. All suits are charging. All radios have been charged to full and removed from the charging docks.

Summary of GreenHab operations: Monitored and watered GreenHab crops and experimental crops. The door was propped open most of the sol due to warmer temperatures.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 5g beans

Summary of Science Dome operations: N/A

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: Placed the observatory and Science Dome items in their correct locations. Retrieved needle nose pliers, wrench, and screwdriver from the RAM, they will be returned to the RAM immediately.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Please advise if you would like me to complete the temporary repair on suit 7’s vent panel, or if you would like me to leave it as-is. The vent panel is secure, but the left side of the panel has about 0.5 inches of undesired movement.

Best Regards,

Mason Kuhn

Crew Engineer

Copyright © The Mars Society. All rights reserved. | Main Site